A company just paid this year's dividend of $3.50 per share on its stock. The dividend is expected to grow at 28 percent per year for two years. Thereafter, the dividend will grow at 4.3 percent per year in perpetuity. If the appropriate discount rate is equal to 12 percent, what is the price of the company's stock today?

A. $74

B. $61

C. $70

D. $67
Answer #1

D1=(3.5*1.28)=4.48

D2=(4.48*1.28)=5.7344

Value after year 2=(D2*Growth rate)/(Discount rate-Growth rate)

=(5.7344*1.043)/(0.12-0.043)

=77.6750545

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.48/1.12+5.7344/1.12^2+77.6750545/1.12^2

=$70(Approx).